OpenStudy (anonymous):

CAN SOMEONE HELP! One diagonal of a rhombus has endpoints (-6, 9) and (-2, 1). What are the endpoints of the other diagonal? (2, 8) and (-10, 2) (-10, 17) and (2, 7) (-6, 5) and (-2, 4) (2, 4) and (6, 6)

OpenStudy (anonymous):

midpoint of the given diagonal is (-4,5) so the midpoint of the given diagonals must be (-4,5) (2, 8) and (-10, 2)

OpenStudy (anonymous):

The mass of an object is equal to the product of the object's density and volume. The density of steel is 7.8 grams per cubic centimeter. What is the volume of a piece of a sheet of steel that has a mass of 3.75 kilograms? 0.4808 cubic centimeters 2.08 cubic centimeters 480.8 cubic centimeters 29.25 cubic centimeters 29,250 cubic centimeters

OpenStudy (anonymous):

0.4808 cubic centimeters

OpenStudy (anonymous):

What is the radius of a sphere with a surface area of 100πcm2? 10 cm 5 cm 25 cm 3 cm

OpenStudy (anonymous):

5 cm

OpenStudy (anonymous):

A circle with the equation (x + 3)2 + (y - 2)2 = 25 is reflected over the line x = 2. What is the equation of the image? (x - 3)2 + (y + 2)2 = 25 (x - 7)2 + (y + 2)2 = 25 (x - 7)2 + (y - 2)2 = 25 (x - 3)2 + (y - 2)2 = 25